Operations Report – November 4th

Crew 214 Operations Report 4-11-2019
SOL: 8
Name of person filing report: Andrew Wheeler
Non-nominal systems: Opportunity, Suit 10
Notes on non-nominal systems: Opportunity brakes do not work, Suit 10 is not functioning

Generator: run
Hours run: 13h 15m
From what time last night: 1845
To what time this morning: 0800
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night): 94%

Diesel Reading – 60%
Station Propane Reading – 82%
Ethanol Free Gasoline: N/A

Water (loft tank): 32 gallons
Water Meter: 1455633 units
Water (static tank): 246 gallons
Static to Loft Pump used – no
Water in Green Hab: 225 gallons
Water in Science Dome: 0 gallons
Toilet tank emptied: yes

Deimos rover used: not at hab
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –

Spirit rover used: yes
Hours: 107.4
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 52%
Currently charging: yes

Opportunity rover used: no
Hours: 65.5
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Curiosity rover used: no
Hours: 113.4
Beginning charge: (Before EVA) : 100%
Ending charge: (On return from EVA, before charging): 100%
Currently charging: yes

Notes on rovers: no

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
ATV Fuel Used: 0 Gals
# Hours the ATVs were used today: 0
Notes on ATVs: Assigned to the director

HabCar used and why, where? N/A

CrewCar used and why, where? N/A

General notes and comments: N/A

Summary of internet: Functional
Summary of suits and radios: 5 Radios and 3 earpieces that need servicing have been identified and set aside, the battery for suit 10 is fully charged and it is Suit 10 that is not functional, further information on Suit 5’s helmet has been forwarded to Capcom.
Summary of Hab operations: All systems nominal
Summary of GreenHab operations: Atmospheric equalization to the green hab was carried out during the day, plants were watered.
Summary of ScienceDome operations: All systems nominal. Sample processing following EVA was carried out.
Summary of RAM operations: Not used
Summary of any observatory issues: N/A
Summary of health and safety issues: All crew healthy
Questions, concerns and requests to Mission Support: Acknowledge that the request for 6 radio batteries is being processed.

Copyright © The Mars Society. All rights reserved. | Main Site